OpenStudy (anonymous):

H(t) = -16t2 + 60t + 82 Factor No direct answers please

OpenStudy (anonymous):

Reduce your given f(x) by taking a common number

OpenStudy (anonymous):

....there is no gcf

OpenStudy (anonymous):

At the very least, as far as constant coefficients are concerned, you can pull out \(-2\). \[-16t^2+60t+82=-2\left(8t^2-30t+41\right)\] Since \(41\) is prime, you know that if the quadratic can be factored, it will be in the form \[-2(ax-41)(bx-1)\] If you expand this you get \[-2(abx^2-(a+41b)x+41)\] which means \(ab=8\) and \(-30=a+41b\).

OpenStudy (anonymous):

Oops should be \(a+41b=\color{red}+30\)